# Hybrid Action Based Reinforcement Learning for Multi-Objective Compatible Autonomous Driving

### v1

@@ III-B Policy and Value Function Representation
$\begin{split}{{\bar{Q}}_{i}}\left({{s_{t}},o,{a_{o,t}}}\right)&=\mathbb{E}{{}_{j\in\left[{1,\cdots,M}\right]}}\left[{{Q_{ij}}\left({{s_{t}},o,{a_{o,t}};\theta_{ij}^{Q}}\right)}\right]\\ &=\frac{1}{M}\sum\limits_{j=1}^{M}{{Q_{ij}}\left({{s_{t}},o,{a_{o,t}};\theta_{ij}^{Q}}\right)\left|{{}_{\forall o\in{\cal O}}}\right.}.\end{split}$ (6)
@@
$\begin{split}{{\cal L}_{t}}\left({{\theta^{\mu}}}\right)=-\frac{1}{M}\sum\limits_{i=1}^{N}{{\omega_{i}}\sum\limits_{j=1}^{M}{\sum\limits_{o\in{\cal O}}{{Q_{ij}}({s_{t}},o,{a_{o,t}};\theta_{ij}^{Q})}}}.\end{split}$ (15)
@@ V-B1 Testing with Rule-Based SVs
For driving efficiency, HPA-MoEC achieves the highest AS through more flexible lane changes. In comparison, SAC-H and PPO-H have lower ASs due to reduced lane-changing flexibility, leading to suboptimal efficiency. Specifically, compared to SAC-H, HPA-MoEC improves AS by 13% and increases NL by 29%. Compared to the above methods with hybrid actions, SAC-C’s direct control of the EV results in the lowest AS and the fewest NL, indicating its inability to effectively leverage lane-changing opportunities to increase speed. Relying on discrete actions, DQN achieves higher AS in some episodes by frequent lane changes, but its overall AS ranks second to last. Overall, the hybrid actions provide greater flexibility and thus improve driving efficiency, especially by using a parameterized action space to generate outputs rather than discretizing part of the continuous actions.

### v2

@@ III-B Policy and Value Function Representation
$\begin{split}{{\bar{Q}}_{i}}\left({{s_{t}},o,{a_{o,t}}}\right)&=\mathbb{E}{{}_{j\in\left[{1,\cdots,M}\right]}}\left[{{Q_{ij}}\left({{s_{t}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q}}\right)}\right]\\ &=\frac{1}{M}\sum\limits_{j=1}^{M}{{Q_{ij}}\left({{s_{t}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q}}\right)\left|{{}_{\forall o\in{\cal O}}}\right.}.\end{split}$ (6)
@@
$\begin{split}{{\cal L}_{t}}\left({{\theta^{\mu}}}\right)=-\frac{1}{M}\sum\limits_{i=1}^{N}{{\omega_{i}}\sum\limits_{j=1}^{M}{\sum\limits_{o\in{\cal O}}{{Q_{ij}}({s_{t}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q})}}}.\end{split}$ (15)
@@ V-B1 Testing with Rule-Based SVs
For driving efficiency, HPA-MoEC achieves the highest AS through more flexible lane changes. In comparison, SAC-H and PPO-H have lower ASs due to reduced lane-changing flexibility, leading to suboptimal efficiency. Specifically, compared to SAC-H, HPA-MoEC improves AS by 13% and increases NL by 28%. Compared to the above methods with hybrid actions, SAC-C’s direct control of the EV results in the lowest AS and the fewest NL, indicating its inability to effectively leverage lane-changing opportunities to increase speed. Relying on discrete actions, DQN achieves higher AS in some episodes by frequent lane changes, but its overall AS ranks second to last. Overall, the hybrid actions provide greater flexibility and thus improve driving efficiency, especially by using a parameterized action space to generate outputs rather than discretizing part of the continuous actions.
@@ VI Conclusion and Future Work
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grant No. 52325212 and No.52372394, in part by the National Key R&D Program of China under Grant No. 2022YFE0117100.